# Cooling and heating regions of Joule-Thomson expansion for AdS black holes: Maxwell-power-Yang-Mills and Kerr Sen black holes

### v1

@@ 6 Conclusion
In this article, we investigated the behavior of two structurally different black holes under the JTE thermodynamic process. The JTE is a process that cools or heats a system by changing its pressure and volume at constant enthalpy, and its main goal is to study the behavior of the $\mu$ coefficient, which describes the change in temperature during the expansion or compression of a system. But before presenting the results, we must first explain the motivations that led us to choose these particular black holes. Recently, Einstein-Power-Young-Mills black hole with AdS structure under JTE process was investigated in a research paper[2, 3]. This raised the question in our mind that if a fields in the Maxwellian form are added to the elements in the action of the above article, can this change and addition of a special form of the field cause a different thermodynamic behavior in their JT representation? Similarly, in the past, the behavior of various forms of rotating black holes was also investigated in this thermodynamic process. But what effect can the addition of Sen-type dilaton fields and antisymmetric tensor fields have on its thermodynamic properties and JT representation? This was a question that could be a good motivation for our investigation. For this purpose, we plotted the isenthalpic curves and the inversion curves for each of the two black holes, for different values of the free parameters, which can be referred to the following results for each. For the AMPYM black hole, the Hawking temperature has zero points that depend on the free parameters and the horizon radius, and the JTC diverges at these points. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole is cooling during the expansion process, and for $P>P_{i}$, the isenthalpic curve has a negative slope, which means that the black hole experiences heating during the expansion process. The inversion curves for different free parameters show that the inversion curve has only one branch and the inversion temperature increases steadily with the inversion pressure, but the slope of the inversion curves becomes smaller. Also, the graphs showed that for low pressure, the inversion temperature varies with the free parameters specified in each graph. Studying the effect of parameters was interesting for us because we observed that the slope of the inversion curve increases with changes in different values of each parameter. But if we want to talk about the effect of Maxwell’s charges as a special result of our work, we must state that, we have found that the ratio of minimum to critical temperature for a 4-dimensional black hole with Yang-Mills hair depends on the values of $q$, $M$, and $C$. For $q>1$, the ratio is almost equal to $1/2$ regardless of the Yang-Mills parameter when "$C\gg 1"$. For $q=1/2$, the ratio is exactly equal to $1/2$ for all values of $C$. For $11/2<q<$ 1, the ratio approaches $1/2$ as $C$ increases. For $20<q<1/$ 2, the ratio deviates from $1/2$ as $C$ increases. These results show that the Maxwell charge can affect the thermodynamic behavior of the black hole and make it more or less similar to a 4-dimensional charged black hole. Also, for the AKS black hole, we found that its parameters, i.e. a and b, can play a vital role in the representation of the value of the ratio of $T_{i}^{min}/T_{C}$, so that by reducing these parameters, the value of the ratio increases. The results are shown in Figure 7 and are summarized in Tables 1 to 5. For the AdS Kerr Sen black, It was very difficult to find the analytical solution for drawing the graphs, so we used numerical solutions to draw them. For $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole is cooling during the expansion process, but with the increases of the black hole parameters, we found that the slope of the figures will change. Also, if we assume the parameter b ( dyonic charge) to be zero, our equations and graphs will simplify to the Kerr-AdS black holes, whose results are thoroughly discussed in[1]. In general, it can be said about this black hole the isenthalpic curves of the AKS black hole show cooling or heating behavior depending on the inversion curve, which is affected by the mass and the parameter $b,a$ of the black hole. The inversion curve has a single branch with a positive slope that decreases with the free parameters, and the inversion temperature and pressure are related by some equations that are numerically solved. The Hawking temperature of the AKS black hole has zero points that depend on the free parameters and the horizon radius and lead the JTC to reflect divergence behavior.
Figure 7: It shows the plot of $(\frac{T_{i}^{min}}{T_{c}})$ in terms of $C$ with respect to free parameters mentioned in the plot.
$q=0.3,n=2,Q_{1}=1$
          $T_{i}^{min}$
          $T_{c}$
          $T_{i}^{min}/T_{c}$
          $C=0.1$
          $0.171532$
          $0.355378$
          $0.482674$
          $C=0.2$
          $0.068796$
          $0.148927$
          $0.461947$
          $C=0.3$
          $0.035473$
          $0.082056$
          $0.432299$
          $C=1$
          Not Exist
          Not Exist
          Not Exist
          $C=5$
          Not Exist
          Not Exist
          Not Exist
          $C=10$
          Not Exist
          Not Exist
          Not Exist
Table 1: Summary of the results for the AMPYM black hole for $q=0.3$
$q=0.5,n=2,Q_{1}=1$
          $T_{i}^{min}$
          $T_{c}$
          $T_{i}^{min}/T_{c}$
          $C=0.1$
          $0.034331$
          $0.068662$
          $0.5$
          $C=0.2$
          $0.017165$
          $0.034330$
          $0.5$
          $C=0.3$
          $0.011443$
          $0.022886$
          $0.5$
          $C=1$
          $0.003433$
          $0.006866$
          $0.5$
          $C=5$
          $0.000686$
          $0.001372$
          $0.5$
          $C=10$
          $0.000343$
          $0.000686$
          $0.5$
Table 2: Summary of the results for the AMPYM black hole for $q=0.5$
$q=0.9,n=2,Q_{1}=1$
          $T_{i}^{min}$
          $T_{c}$
          $T_{i}^{min}/T_{c}$
          $C=0.1$
          $0.019146$
          $0.03874$
          $0.493657$
          $C=0.2$
          $0.018873$
          $0.038208$
          $0493962$
          $C=0.3$
          $0.018446$
          $0.037309$
          $0.494405$
          $C=1$
          $0.013846$
          $0.027832$
          $0497502$
          $C=5$
          $0.004141$
          $0.008286$
          $0.499762$
          $C=10$
          $0.002133$
          $0.004267$
          $0.499921$
Table 3: Summary of the results for the AMPYM black hole for $q=0.9$
          ($a,b$)
          $T_{i}^{min}$
          $T_{c}$
          $T_{i}^{min}/T_{c}$
($0.0099$, $0.00499$)
          $0.0787687$
          $0.335030$
          $0.235109$
($0.00972$, $0.00486$)
          $0.0782007$
          $0.2302899$
          $0.33957$
($0.00951$, $0.00475$)
          $0.078821$
          $0.156982$
          $0.5$
($0.00937$, $0.00468$)
          $0.0788340$
          $0.116006$
          $0.67379$
($0.00892$, $0.00442$)
          $0.776375$
          $0.0454247$
          $1.709$
@@
       AMPYM BH
          $q>1$, $C\gg 1$
          $T_{min}/T_{c}=\frac{1}{2}$
          This paper
       AMPYM BH
$q=0.5$, all values of "$C"$
          $T_{min}/T_{c}\simeq\frac{1}{2}$
          This paper
       AMPYM BH
          $0.5<q\leq 1$ , $C\gg 1$
         $T_{min}/T_{c}\simeq\frac{1}{2}$
          This paper
       AKS BH
          ($a=0.00951$, $b=0.00475$)
          $T_{min}/T_{c}\simeq\frac{1}{2}$
          This paper
Van der Waals fluid
          Exist
          $T_{min}/T_{c}=\frac{3}{4}$
          [24]
          RN-AdS BH
          Exist
          $T_{min}/T_{c}=\frac{1}{2}$
          [25]
d-dimensional AdS BH
          Exist
          $T_{min}/T_{c}<\frac{1}{2}$
          [79]
Gauss-Bonnet BH
          Exist
$T_{min}/T_{c}=0.4765$
          [80]
          torus-like BH
          Not Exist
$T_{min}/T_{c}$= Not Exist
          [81]
          BTZ BH
          Not Exist
$T_{min}/T_{c}$= Not Exist
          [82]
@@
Table 5: Summary of the results for the AMPYM and AKS black holes compare with other results

### v2

@@ 6 Conclusion
In this article, we investigated the behavior of two structurally different black holes under the JTE thermodynamic process. The JTE is a process that cools or heats a system by changing its pressure and volume at constant enthalpy, and its main goal is to study the behavior of the $\mu$ coefficient, which describes the change in temperature during the expansion or compression of a system. But before presenting the results, we must first explain the motivations that led us to choose these particular black holes. Recently, an EPYM black hole with AdS structure under the JTE process was investigated in a research paper[27, 28]. This raised the question in our mind if fields in the Maxwellian form are added to the elements in the action of the above article, can this change and addition of a special form of the field cause a different thermodynamic behavior in their JT representation? Similarly, in the past, the behavior of various forms of rotating black holes was also investigated in this thermodynamic process. But what effect can the addition of Sen-type dilaton fields and antisymmetric tensor fields have on its thermodynamic properties and JT representation? This was a question that could be a good motivation for our investigation. For this purpose, we plotted the isenthalpic curves and the inversion curves for each of the two black holes, for different values of the free parameters, which can be referred to the following results for each. For the AEMPYM black hole, the Hawking temperature has zero points that depend on the free parameters and the horizon radius, and the JTC diverges at these points. The isenthalpic curves are divided into two regions by the inversion curve: for $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole is cooling during the expansion process, and for $P>P_{i}$, the isenthalpic curve has a negative slope, which means that the black hole experiences heating during the expansion process. The inversion curves for different free parameters show that the inversion curve has only one branch and the inversion temperature increases steadily with the inversion pressure, but the slope of the inversion curves becomes smaller. Also, the graphs showed that for low pressure, the inversion temperature varies with the free parameters specified in each graph. Studying the effect of parameters was interesting for us because we observed that the slope of the inversion curve increases with changes in the different values of each parameter. But if we want to discuss the effect of Maxwell’s charges as a special result of our work, we must state that, we have found that the ratio of minimum to critical temperature for a 4-dimensional black hole with Yang-Mills hair depends on the values of $q$, $M$, and $C$. For $q=1/2$, the ratio is exactly equal to $1/2$ for all values of $C$. For $11/2<q<$ 1, the ratio approaches $1/2$ as $C$ increases. For $20<q<1/$ 2, the ratio deviates from $1/2$ as $C$ increases. These results show that the Maxwell charge can affect the thermodynamic behavior of the black hole and make it more or less similar to a 4-dimensional charged black hole. Also, for the AKS black hole, we found that its parameters, i.e. a and b, can play a vital role in the representation of the value of the ratio of $T_{i}^{min}/T_{C}$ so that by reducing these parameters, the value of the ratio increases. The results are shown in Fig. (4) and are summarized in tables (1) to (5). For the AKS black, It was very difficult to find the analytical solution for drawing the graphs, so we used numerical solutions to draw them. For $P<P_{i}$, the isenthalpic curve has a positive slope, indicating that the black hole is cooling during the expansion process, but with the increases of the black hole parameters, we found that the slope of the figures will change. Also, if we assume the parameter b ( dyonic charge) to be zero, our equations and graphs will simplify to the Kerr-AdS black holes, whose results are thoroughly discussed in[26]. In general, it can be said about this black hole the isenthalpic curves of the AKS black hole show cooling or heating behavior depending on the inversion curve, which is affected by the mass and the parameter $b,a$ of the black hole. The inversion curve has a single branch with a positive slope that decreases with the free parameters, and the inversion temperature and pressure are related by some equations that are numerically solved. The Hawking temperature of the AKS black hole has zero points that depend on the free parameters and the horizon radius and lead the JTC to reflect divergence behavior.
$q=0.3,n=2,Q=1$           $T_{i}^{min}$           $T_{c}$           $T_{i}^{min}/T_{c}$
          $C=0.1$           $0.171532$           $0.355378$           $0.482674$
          $C=0.2$           $0.068796$           $0.148927$           $0.461947$
          $C=0.3$           $0.035473$           $0.082056$           $0.432299$
          $C=1$           Not Exist           Not Exist           Not Exist
          $C=5$           Not Exist           Not Exist           Not Exist
          $C=10$           Not Exist           Not Exist           Not Exist
Table 1: Summary of the results for the AEMPYM black hole for $q=0.3$
$q=0.5,n=2,Q=1$           $T_{i}^{min}$           $T_{c}$           $T_{i}^{min}/T_{c}$
          $C=0.1$           $0.034331$           $0.068662$           $0.5$
          $C=0.2$           $0.017165$           $0.034330$           $0.5$
          $C=0.3$           $0.011443$           $0.022886$           $0.5$
          $C=1$           $0.003433$           $0.006866$           $0.5$
          $C=5$           $0.000686$           $0.001372$           $0.5$
          $C=10$           $0.000343$           $0.000686$           $0.5$
Table 2: Summary of the results for the AEMPYM black hole for $q=0.5$
$q=0.9,n=2,Q=1$           $T_{i}^{min}$           $T_{c}$           $T_{i}^{min}/T_{c}$
          $C=0.1$           $0.019146$           $0.03874$           $0.493657$
          $C=0.2$           $0.018873$           $0.038208$           $0493962$
          $C=0.3$           $0.018446$           $0.037309$           $0.494405$
          $C=1$           $0.013846$           $0.027832$           $0497502$
          $C=5$           $0.004141$           $0.008286$           $0.499762$
          $C=10$           $0.002133$           $0.004267$           $0.499921$
Table 3: Summary of the results for the AEMPYM black hole for $q=0.9$
          ($a,b$)           $T_{i}^{min}$           $T_{c}$           $T_{i}^{min}/T_{c}$
($0.0099$, $0.00499$)           $0.0787687$           $0.335030$           $0.235109$
($0.00972$, $0.00486$)           $0.0782007$           $0.2302899$           $0.33957$
($0.00951$, $0.00475$)           $0.078821$           $0.156982$           $0.5$
($0.00937$, $0.00468$)           $0.0788340$           $0.116006$           $0.67379$
($0.00892$, $0.00442$)           $0.776375$           $0.0454247$           $1.709$
@@
       AEMPYM BH           $q>1$, $C\gg 1$           $T_{min}/T_{c}=\frac{1}{2}$           This paper
       AEMPYM BH $q=0.5$, all values of "$C"$           $T_{min}/T_{c}\simeq\frac{1}{2}$           This paper
       AEMPYM BH           $0.5<q\leq 1$ , $C\gg 1$          $T_{min}/T_{c}\simeq\frac{1}{2}$           This paper
       AKS BH           ($a=0.00951$, $b=0.00475$)           $T_{min}/T_{c}\simeq\frac{1}{2}$           This paper
Van der Waals fluid           Exist           $T_{min}/T_{c}=\frac{3}{4}$           [23]
          RN-AdS BH           Exist           $T_{min}/T_{c}=\frac{1}{2}$           [29]
d-dimensional AdS BH           Exist           $T_{min}/T_{c}<\frac{1}{2}$           [110]
Gauss-Bonnet BH           Exist $T_{min}/T_{c}=0.4765$           [111]
          torus-like BH           Not Exist $T_{min}/T_{c}$= Not Exist           [112]
          BTZ BH           Not Exist $T_{min}/T_{c}$= Not Exist           [113]
@@
Table 5: Summary of the results for the AEMPYM and AKS black holes compare with other results
7 Appendix: Numerical technique
To construct the temperature-pressure graph $(T(M,P))$ as depicted in Figure (2), it is essential to first determine the radius of the event horizon $(r_{+})$ as a function of pressure $(P)$, utilizing Eq.(5.1). Subsequently, by substituting this relationship into Eq.(17), we can generate the desired graph. However, the intricate nature of Eq.(5.1) precludes us from deriving an analytical expression for the event horizon radius $(r_{+})$ in terms of pressure $(P)$. As a result, we resort to numerical methods for this purpose. By holding the mass $(M)$, the charge $(Q)$, and the constant $(C)$ fixed, and by inputting various radii for the event horizon into Eqs. (5.1) and (5.2), we can calculate the corresponding temperatures and pressures for different events. This numerical approach applies to any scenario that defies analytical solutions. Indeed, this method can be employed for both types of black holes as delineated by the equations presented in the text. To illustrate this numerical technique, we provide an example below, which corresponds to the scenario outlined in Fig. (2). Please note that the numerical approach for Fig. (2) will be placed in the following table (6) for the AEMPYM black hole with respect to its free parameters demonstrating the numerical method to derive the temperature and pressure values associated with the event horizon of the black hole. Also, to draw Fig. (5) for the AKS black hole, a similar process can be done according to Eqs. (5.12) and (5.14), as well as its free parameters in the text.
                       P                        T
                 0.196046                  0.011344
                 0.193908                  0.0822105
                 0.189191                  0.137814
                 0.182693                  0.181027
                 0.17897                   0.198729
                 0.170874                  0.227563
                 0.162223                  0.249009
                 0.153304                  0.264447
                 0.144326                  0.275003
                 0.135443                  0.281598
                 0.131072                  0.283651
                 0.122525                  0.285673
                 0.119611                  0.285807
                 0.118364                  0.285783
                 0.106396                  0.283216
                 0.0883264                  0.272013
                 0.0592559                  0.236967
                 0.0382974                  0.198038
                 0.0234604                  0.161811
                 0.0130321                  0.130287
                 0.00572774                  0.103673
                 0.000628043                  0.0815221
                 0.000212542                  0.0795286
Table 6: $M=4,q=1.5,n=3,C=0.5,Q=1$
To clarify the relationship between the critical temperature, $(T_{c})$, and Maxwell’s charge, $(C)$, we analyzed the (AEMPYM) black holes for different values $Q=1,n=2,q=2$ and $Q=1,n=2,q=1.5$. Using the critical condition, we first obtain $r_{+,c}$. Then by placing it in the critical temperature, we calculate the direct relationship between the critical temperature and the Maxwell charge. Also, to better understand this relationship, we draw Fig. (8). As shown in Fig. (8), we find for larger values of $(C)$, the results tend to converge, becoming consistent across different $(q)$ values. Conversely, for Maxwell charge values less than 1, a higher critical temperature is noted for larger $(q)$ values.
[…]
where for $Q=1,n=2,q=2$, we will have,
$\begin{split}r_{+,c}=0.5\sqrt{2\sqrt{36c^{4}+84.8528}+12c^{2}}\end{split}$ (7.2)
Then, we can obtain,
[…]
Also, for $Q=1,n=2,q=1.5$, we will have,
$\begin{split}r_{+,c}=0.5\sqrt{2\sqrt{36c^{4}+84.8528}+12c^{2}}\end{split}$ (7.4)
In this case, the corresponding critical temperature is obtained from the following equation,
$\begin{split}&\mathcal{Y}_{1}=162.975\bigg{[}0.00195313\bigg{(}2\sqrt{36c^{4}+84.8528}+12c^{2}\bigg{)}^{5}-0.015625c^{2}\bigg{(}2\sqrt{36c^{4}+84.8528}+12c^{2}\bigg{)}^{4}\\ &-1.06066\bigg{(}1\sqrt{36c^{4}+84.8528}+6c^{2}\bigg{)}^{3}\bigg{]}\\ &\mathcal{Y}_{2}=(2\sqrt{36c^{4}+84.8528}+12c^{2})^{5.5}\\ &T_{c}=\frac{\mathcal{Y}_{1}}{\mathcal{Y}_{2}}\end{split}$ (7.5)
Figure 8: The plot of $T_{c}-C$. The solid blue line illustrated with $q=2,n=2$, and $Q=1$ and solid red line denoted with $q=1.5,n=2$, and $Q=1$ for the AEMPYM black hole